OpenStudy (anonymous):

A test has 20 questions. If Peter gets 80 percent correct, how many questions did Peter miss? A. 2 B. 6 C. 5 D. 4

jimthompson5910 (jim_thompson5910):

80% of 20 = 0.80*20 = ???

OpenStudy (anonymous):

HOWCAN U EXPLAIN THE ANSWER I KNOW ITS d

jimthompson5910 (jim_thompson5910):

what's 80% of 20

OpenStudy (anonymous):

YEAH WAT IS IT

jimthompson5910 (jim_thompson5910):

0.80*20 = ???

jimthompson5910 (jim_thompson5910):

use a calculator if you need to

OpenStudy (anonymous):

ITS GIVES 16

jimthompson5910 (jim_thompson5910):

so he got 16 correct

jimthompson5910 (jim_thompson5910):

and got 20 - 16 = 4 wrong

OpenStudy (anonymous):

RIGHT THANKSSS BERLYY STARTING

jimthompson5910 (jim_thompson5910):

you're welcome